mu ud fund compan u ers ts cus urmers d variety o funds: d mune -marke und the parcantagas af custamars in tha differant funds are as fallows. rerent buid unds shur intermediate aru on und rete dnu hig 1-tisk dnd d d dnced und in ng customers wh u nn shares in us one fun ree errn v o s o ma u Maney-market Short hond Intermediate bond 5% Long bond 25% 14% High-risk stock Moderate-risk stock Balanced 2596 A customer who ouns shares in just one tund is randomly selected. (a) what is the probability that the selected individual owns shares in the balanced fund? What is the probability that the individual owns shares in a bond tund? What is the probability that the selected individual does not own shares in a stock fund?
Answer #1

a) P(Owns share in balanced fund) = 0.10

b) P(Owns share in a bond fund) = 0.14 + 0.05 + 0.05 = 0.24

c) P(Doesn't own share in a stock fund) = 1 - (0.16 + 0.25) = 0.59